OpenStudy (loser66):

The medians of a right triangle which are drawn from the vertices of the acute angles are 5 and \(\sqrt{40}\). what is the length of the hypotenuse? Please, help

OpenStudy (ikram002p):



OpenStudy (ikram002p):

\((2x)^2+y^2= 40\\ x^2+(2y)^2=25\)

OpenStudy (ikram002p):

so you wanna h \(h=\sqrt{4x^2+4y^2 }=\sqrt{4(9+4)}=\sqrt{52}\)

OpenStudy (loser66):

Thank you so much, my friend.

OpenStudy (ikram002p):

you are welcome anytime, my friend.